OpenStudy (anonymous):

What is the slope of the line through the points (4,−7) and (7,−9)?

OpenStudy (igreen):

Plug them in the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-9+7}{7-4}\) Add & Subtract those, then divide, what do you get?

OpenStudy (igreen):

@Maddox1000

OpenStudy (anonymous):

I know, It's just hard is all.

OpenStudy (anonymous):

Dangit, I don't know how to do this. It was so much easier in the pretests.